40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 27 02:40:06 2009, pete, Summary, IOO, MC glitch investigation MC1_Drift.png
    Reply  Fri Mar 27 15:05:42 2009, Yoichi, Update, IOO, MC glitch investigation MC1_Drift.pdfMC2_Drift.pdf
       Reply  Fri Mar 27 17:52:16 2009, Yoichi, Update, IOO, MC glitch investigation 
Message ID: 1435     Entry time: Fri Mar 27 02:40:06 2009     Reply to this: 1437
Author: pete 
Type: Summary 
Category: IOO 
Subject: MC glitch investigation 

Yoichi, Pete

The MC loses lock due to glitches in the MC1 coils. 
We do not know which coil for sure, and we do not know if it is a problem going into the board, or a problem on the board. 
We suspect either the UL or LR coil bias circuits (Pete would bet on UL).  If you look at the bottom 4 plots in the attached file, you can see a relatively large 3 minute dip in the UL OSEM output, with a corresponding bump in the LR (and smaller dips in the other diagonal).  
These bumps do not show up in the VMONS which is why we are suspicious of the bias.
To test we are monitoring 4 points in test channels, for UL and UR, both going into the bias driver circuit, and coming out of the current buffer before going into the coils. 
 

We ran cable from the suspension rack to the IOO rack to record the signals with DAQ channels.

The test channels:

UL coil      C1:IOO-MC_DRUM1  (Caryn was using, we will replace when we are done)

UL input   C1:IOO-MC_TMP1 (Caryn was using, we will replace when we are done)

LR coil      C1:PEM-OSA_SPTEMP

LR input   C1:PEM-OSA_APTEMP

We will leave these overnight; we intend to remove them tomorrow or Monday.

We closed the PSL shutter and killed the MC autolocker.

Attachment 1: MC1_Drift.png  36 kB  | Hide | Hide all
MC1_Drift.png
ELOG V3.1.3-